### v1

# Semitoric Families on Pentagon Spaces

Abstract.
Semitoric systems are a special type of 4-dimensional integrable system where one of the functions is the moment map of a Hamiltonian $S^{1}$-action. These systems were classified by Pelayo and Vũ Ngọc, but there are not many explicit examples. Recently, Le Floch and Palmer introduced semitoric transition families in which a singular point transitions between elliptic-elliptic type and focus-focus type as the parameter varies. In this paper, we construct semitoric transition families on pentagon spaces by interpolating two semitoric systems of toric type. More specifically, we show that there exists a unique singular point that changes from elliptic-elliptic type to focus-focus type and back to elliptic-elliptic type as the parameter varies.
2020 Mathematics Subject Classification:
37J35, 53D20
1. Introduction
Let $(M,\omega)$ be a symplectic manifold. Unless otherwise specified, we always assume that $M$ is connected. Recall that to each smooth function $f:M\to{\mathbb{R}}$, one can associate its Hamiltonian vector field $X_{f}$ given by the equation $df=\omega(X_{f},\cdot)$. Given two functions $f,g:M\to{\mathbb{R}}$, their Poisson bracket $\{f,g\}$ is another function defined by $\{f,g\}:=\omega(X_{f},X_{g})$. We say two functions $f,g$ Poisson commute if $\{f,g\}=0$.
An integrable system on a $2n$-dimensional symplectic manifold $(M,\omega)$ is a map $F=(f_{1},\ldots,f_{n}):M\to{\mathbb{R}}^{n}$ such that $\{f_{i},f_{j}\}=0$ for all $i,j$ and $df_{1},\ldots,df_{n}$ are linearly independent almost everywhere. A semitoric system is a special kind of integrable system when $\dim M=4$. More specifically, an integrable system $(M,\omega,F=(J,H))$ is a semitoric system if $J$ is a proper map that generates an effective Hamiltonian $S^{1}$-action on $M$ and the momentum map $F$ only has non-degenerate singularities without hyperbolic type (see Definition 3.3). Semitoric systems were introduced by  [VuN07] following the work of [Sym03]. Pelayo and Vũ Ngọc [PVN09, PVN11] first classified semitoric systems with generic critical fibers. Following their work, Pelayo-Palmer-Tang [PPT24] generalized the classification to any semitoric systems. Semitoric systems have been intensively studied in the recent years. See [ADH20, ADH19, AHP25, LFP19] for studies of invariants of semitoric systems and [AH19, SVN17, DKL+23] for recent survey articles and additional references.
The coupled angular momenta system was the first example of a family of semitoric system known to us. It was first introduced in [SZ99] and studied in detail in [LFP19]. More families were studied later by many people; see [HP18, DP16, DMH21, LFP24, LFP19, LFP23]. In [LFP24], the authors gave the definition of semitorc families with degenerate times and semitoric transition families. A semitoric family with degenerate times $t_{1},\dots,t_{n}\in[0,1]$, is a family of integrable systems $(M,\omega,F_{t})$, $0\leq t\leq 1$ on a 4-dimensional symplectic manifold $(M,\omega)$ such that $(M,\omega,F_{t}:=(J,H_{t}))$, where $H_{t}$ is of the form $H_{t}=H(t,\cdot)$ for some smooth $H:[0,1]\times M\rightarrow{\mathbb{R}}$, is semitoric if and only if $t\in[0,1]\setminus\{t_{1},\dots,t_{n}\}$. In this paper, we study semitoric transition families on pentagon spaces.
A semitoric transition family with transition point $p\in M$ and transition times $t^{-},t^{+}$, with $0<t^{-}<t^{+}<1$, is a family of integrable systems $(M,\omega,F_{t}=(J,H_{t}))$, $0\leq t\leq 1$ on a 4-dimensional symplectic manifold $(M,\omega)$ such that:
  • $(M,\omega,F_{t})$ is semitoric for $t\in[0,1]\setminus\{t^{-},t^{+}\}$;
  • for $t\in[0,t^{-})\cup(t^{+},1]$, the singular point $p$ is of elliptic-elliptic type;
  • for $t\in(t^{-},t^{+})$, the singular point $p$ is of focus-focus type;
  • for $t=t^{-}$ and $t=t^{+}$, there are no degenerate singular point in $M\setminus\{p\}$, and
  • if $p$ is a maximum (respectively minimum) of $(H_{0})|_{J^{-1}(J(p))}$ then $p$ is a minimum (respectively maximum) of $(H_{1})|_{J^{-1}(J(p))}$.
Let $r_{1},\ldots,r_{n}\in{\mathbb{R}}_{+}$ be positive numbers. Define the space of $n$-gons with side lengths $r_{1},\ldots,r_{n}$ to be the set
$\widetilde{\mathcal{P}}(r_{1},\ldots,r_{n}):=\{\rho:\{1,\ldots,n\}\to{\mathbb{R}}^{3}\mid\sum_{i=1}^{n}\rho_{i}=0,|\rho_{i}|=r_{i}\}.$
The $\mathbf{n}$-polygon space $\mathcal{P}(r_{1},\ldots,r_{n})$ is the quotient space $\mathcal{P}(r_{1},\ldots,r_{n}):=\widetilde{\mathcal{P}}(r_{1},\ldots,r_{n})/SO(3)$. In other words, the polygon space consists of the configurations in ${\mathbb{R}}^{3}$ of polygons with side lengths $r_{1},\ldots,r_{n}$, up to rotation. We say $r_{1},\ldots,r_{n}$ are generic if there does not exist a map $\epsilon:\{1,\ldots,n\}\to\{\pm 1\}$ such that $\sum_{i=1}^{n}\epsilon_{i}r_{i}=0$. Generic side lengths guarantee that the polygon space, if not empty, is a smooth manifold of dimension 2n62n-6. These spaces come with symplectic structures. Consider $\widetilde{M}:=S^{2}_{r_{1}}\times\cdots\times S^{2}_{r_{n}}$, the product of spheres of radii $r_{1},\ldots,r_{n}$, with symplectic form $\frac{1}{r_{1}^{2}}\omega\oplus\cdots\oplus\frac{1}{r_{n}^{2}}\omega$, where $\omega_{u}(v,w)=u\cdot(v\times w)$. Here, we think of $S^{2}_{r_{i}}$ as a submanifold of ${\mathbb{R}}^{3}$. The map $\mu:\widetilde{M}\to{\mathbb{R}}^{3}$ given by $\mu(\rho_{1},\ldots,\rho_{n})=\sum_{i=1}^{n}\rho_{i}$ is the moment map of the diagonal Hamiltonian $SO(3)$-action on $\widetilde{M}$, where we identify $\mathfrak{so}_{3}^{*}$ with ${\mathbb{R}}^{3}$ equipped with the cross product. Then the symplectic reduction at level 0 gives us $\mathcal{P}(r_{1},\ldots,r_{n})=\mu^{-1}(0)/SO(3)$, so it comes with a symplectic structure. See [HK97, Kly94, KM95, KM96]. Let $I\subset\{1,\ldots,n\}$ be a nonempty proper subset. Define $\ell_{I}:\mathcal{P}(r_{1},\ldots,r_{n})\to{\mathbb{R}}$ by $\ell_{I}([\rho])=|\sum_{i\in I}\rho_{i}|$.
Theorem 1.1.
Let $r_{1},\ldots,r_{5}$ be positive numbers such that
  • $\frac{1}{2}r_{5}\leq r_{4}<r_{5}<r_{3}$,
  • $|r_{1}+r_{2}-r_{3}|<r_{5}-r_{4}$, and
  • $|r_{3}-r_{4}-r_{5}|<|r_{1}-r_{2}|<r_{3}+r_{4}-r_{5}$.
Then the family $F_{t}:=(\ell_{12},t\ell_{34}^{2}+(1-t)\ell_{45}^{2})$ is a semitoric transition family on $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$, with transition point $P$ represented by the planar pentagon, whose five edges are given by
$\rho_{1}=\left(\frac{r_{1}^{2}-r^{2}_{2}+j^{2}}{2j},\frac{\sqrt{\big((r_{1}+j)^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j)^{2}\big)}}{2j},0\right),$
$\rho_{2}=\left(\frac{r_{2}^{2}-r_{1}^{2}+j^{2}}{2j},-\frac{\sqrt{\big((r_{1}+j)^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j)^{2}\big)}}{2j},0\right),$
$\rho_{3}=(-r_{3},0,0),\rho_{4}=(-r_{4},0,0),\rho_{5}=(r_{5},0,0),$
and transition times
$t^{\pm}=\frac{\,r_{3}r_{4}+r_{3}r_{5}+2r_{4}^{2}-3r_{4}r_{5}+r_{5}^{2}\,\pm\,2\sqrt{r_{3}r_{4}r_{5}j}}{\,(r_{3}+r_{5})^{2}+4r_{4}j\,}.$
Here $j=r_{3}+r_{4}-r_{5}$.
Theorem 1.1 gives new explicit examples of semitoric systems. The index sets in the statement are chosen for convenience. In fact, one may take any three distinct index sets $I,J,K$ of size $2$ such that $I$ is disjoint from both $J,K$ and consider the families $(\ell_{I},t\ell_{J}^{2}+(1-t)\ell_{K}^{2})$ on pentagon spaces whose side lengths satisfy similar constraints to those in Theorem 1.1. Moreover, different choices of side lengths $r_{1},\ldots,r_{5}$ that satisfy the assumptions of Theorem 1.1 give different symplectic forms on pentagon spaces of the same diffeotype, so these families will be useful in understanding how the change of symplectic form affects the transition time and other invariants of these semitoric systems. Additionally, our construction can be easily generalized to polygon spaces of higher dimensions to study integrable systems where some components of the functions give a moment map of a Hamiltonian torus action.
Structure of the paper:
In Section 2, we review some facts about polygon spaces and study an $S^{1}$-action on the pentagon space of interest. In Section 3, we review the basic concepts of semitoric systems and study the local theory of the transition point. In Section 4, we prove that the singular points other than the transition point $P$ are non-degenerate. In Section 5, we study the transition point $P$ and transition times. In Section 6 prove Theorem 1.1.
Acknowledgements.
We would like to express sincere gratitude to Yohann Le Floch, Joseph Palmer, and Susan Tolman for their patient guidance, insightful discussions, and invaluable feedback during the development of this project. The first author also wants to thank Tara Holm for introducing polygon spaces during his visit to Cornell University, and Allen Knutson for some helpful discussions about pentagon spaces. The first author was partially supported by NSF Grant 2204359.
2. Polygon spaces
In this section, we review some properties of polygon spaces. For a nonempty proper subset $I\subsetneq\{1,\ldots,n\}$, we define $\ell_{I}:\mathcal{P}(r_{1},\ldots,r_{n})\to{\mathbb{R}}$ by $\ell_{I}([\rho])=|\sum_{i\in I}\rho_{i}|$. Its Hamiltonian flow is called the bending flow associated to $I$. Whenever $\ell_{I}$ never vanishes on $\mathcal{P}(r_{1},\ldots,r_{n})$, the bending flow associated to $I$ is periodic (see [Kly94, section 2.1] or  [KM96, Corollary 3.9]). It rotates the set of vectors $\{\rho_{i}:i\in I\}$ around the axis $\sum_{i\in I}\rho_{i}$. Hence, $\ell_{I}$ is a moment map of a Hamiltonian $S^{1}$-action on $\mathcal{P}(r_{1},\ldots,r_{n})$. We first show that reducing by these $S^{1}$-actions yields another polygon space. For simplicity, we will pick $I=\{1,2\}$.
Lemma 2.1.
Fix positive numbers $r_{1}>r_{2}$. Let $M:=S^{2}_{r_{1}}\times S^{2}_{r_{2}}$ be equipped with the symplectic form $\frac{1}{r_{1}^{2}}\omega\oplus\frac{1}{r_{2}^{2}}\omega$, where $\omega_{u}(v,w)=u\cdot(v\times w)$. Consider the Hamiltonian $S^{1}$-action on $M$ with moment map $f(\rho_{1},\rho_{2})=|\rho_{1}+\rho_{2}|$. For any $c\in[r_{1}-r_{2},r_{1}+r_{2}]$, $f^{-1}(c)/S^{1}$ is a symplectic manifold symplectomorphic to $(S^{2}_{c},\frac{1}{c^{2}}\omega)$.
Proof.
The Hamiltonian vector field of $f$ is given by $X_{f}(\rho_{1},\rho_{2})=\frac{1}{|\rho_{1}+\rho_{2}|}(\rho_{2}\times\rho_{1},\rho_{1}\times\rho_{2})$. Hence, if $c\in(r_{1}-r_{2},r_{1}+r_{2})$, then $c$ is a regular value of $f$. Moreover, for any $(\rho_{1},\rho_{2})\in f^{-1}(c)$, $S^{1}$ acts on it by rotating it around $\rho_{1}+\rho_{2}$ (see [KM96, Proposition 3.11] for example), so $S^{1}$ acts freely on $f^{-1}(c)$ whenever $c$ is a regular value. It follows that $f^{-1}(c)/S^{1}$ is a symplectic manifold.
Consider the map $\Phi:f^{-1}(c)\to S^{2}_{c}$ that maps $(\rho_{1},\rho_{2})$ to $\rho_{1}+\rho_{2}$. Since $c\in(r_{1}-r_{2},r_{1}+r_{2})$, this map is surjective. Moreover, by the law of cosine, for any $\rho\in S^{2}_{c}$, $\Phi^{-1}(\rho)$ is a single $S^{1}$-orbit. We prove that $\Phi^{*}(\frac{1}{c^{2}}\omega)=\iota^{*}(\frac{1}{r_{1}^{2}}\omega\oplus\frac{1}{r_{2}^{2}}\omega)$, where $\iota:f^{-1}(c)\hookrightarrow M$ is the inclusion. Let $e_{1},e_{2},e_{3}$ be the standard basis of ${\mathbb{R}}^{3}$ and define $V_{i}:=(\rho_{1}\times e_{i},\rho_{2}\times e_{i})$ for $i=1,2,3$. It is straightforward to check that $V_{1},V_{2},V_{3}$ form a basis of $T_{(\rho_{1},\rho_{2})}f^{-1}(c)$.
$(\Phi^{*}\frac{1}{c^{2}}\omega)_{(\rho_{1},\rho_{2})}(V_{i},V_{j})$ $=\frac{\rho_{1}+\rho_{2}}{c^{2}}\cdot[((\rho_{1}+\rho_{2})\times e_{i})\times((\rho_{1}+\rho_{2})\times e_{j})]$
$=\frac{|\rho_{1}+\rho_{2}|^{2}}{c^{2}}(\rho_{1}+\rho_{2})\cdot(e_{i}\times e_{j})$
$=(\rho_{1}+\rho_{2})\cdot(e_{i}\times e_{j}).$
On the other hand,
$\iota^{*}\left(\frac{1}{r_{1}^{2}}\omega\oplus\frac{1}{r_{2}^{2}}\omega\right)_{(\rho_{1},\rho_{2})}(V_{i},V_{j})$ $=\frac{\rho_{1}}{r_{1}^{2}}\cdot[(\rho_{1}\times e_{i})\times(\rho_{1}\times e_{j})]+\frac{\rho_{2}}{r_{2}^{2}}\cdot[(\rho_{2}\times e_{i})\times(\rho_{2}\times e_{j})]$
$=\frac{|\rho_{1}|^{2}}{r_{1}^{2}}\rho_{1}\cdot(e_{i}\times e_{j})+\frac{|\rho_{2}|^{2}}{r_{2}^{2}}\rho_{2}\cdot(e_{i}\times e_{j})$
$=(\rho_{1}+\rho_{2})\cdot(e_{i}\times e_{j}).$
Therefore, by the symplectic reduction Theorem [Mey73, MW74], when $c$ is a regular value, $f^{-1}(c)/S^{1}$ is symplectomorphic to $(S^{2}_{c},\frac{1}{c^{2}}\omega)$.
Now assume that $c=r_{1}\pm r_{2}$. In this case, $X_{f}(\rho_{1},\rho_{2})=0$ for any $(\rho_{1},\rho_{2})\in f^{-1}(c)$, so every point in $f^{-1}(c)$ is a fixed point. It follows that $f^{-1}(c)/S^{1}=f^{-1}(c)$. The map $\Psi:S^{2}_{c}\to f^{-1}(c)$ sending $\rho$ to $(\frac{r_{1}}{c}\rho,\pm\frac{r_{2}}{c}\rho)$ is the inverse of the map $\Phi$ defined above. Moreover, for any $v,w\in T_{\rho}S^{2}_{c}$, we have that
$\Psi^{*}\left(\frac{1}{r_{1}^{2}}\omega\oplus\frac{1}{r_{2}^{2}}\omega\right)_{\rho}(v,w)$ $=\frac{1}{r_{1}^{2}}\frac{r_{1}}{c}\rho\cdot\bigg(\frac{r_{1}}{c}v\times\frac{r_{1}}{c}w\bigg)+\frac{1}{r_{2}^{2}}\frac{\pm r_{2}}{c}\rho\cdot\bigg(\frac{\pm r_{2}}{c}v\times\frac{\pm r_{2}}{c}w\bigg)$
$=\frac{r_{1}}{c^{3}}\rho\cdot(v\times w)\pm\frac{r_{2}}{c^{3}}\rho\cdot(v\times w)$
$=\frac{r_{1}\pm r_{2}}{c^{3}}\rho\cdot(v\times w)$
$=\frac{1}{c^{2}}\rho\cdot(v\times w)$
$=\frac{1}{c^{2}}\omega_{\rho}(v,w)$
Hence, $\Psi$ is indeed a symplectomorphism. The claim follows. ∎
Proposition 2.2.
Fix positive numbers $r_{1},\ldots,r_{n}$ such that $r_{1}\neq r_{2}$ and $\mathcal{P}(r_{1},r_{2},\ldots,r_{n})$ is a smooth manifold. Fix $|r_{1}-r_{2}|\leq c\leq r_{1}+r_{2}$ such that the reduced space $\ell_{12}^{-1}(c)/S^{1}$ is a smooth manifold. Then $\ell_{12}^{-1}(c)/S^{1}$ is symplectomorphic to $\mathcal{P}(c,r_{3},r_{4},\ldots,r_{n})$
Proof.
Let $\widetilde{M}:=S^{2}_{r_{1}}\times\cdots\times S^{2}_{r_{n}}$ be the product of symplectic $2$-spheres of radius $r_{1},\ldots,r_{n}$ and let $\mu:\widetilde{M}\to{\mathbb{R}}^{3}$ be the moment map of the diagonal Hamiltonian $SO(3)$-action on $\widetilde{M}$. Under this setting, $\mathcal{P}(r_{1},\ldots,r_{n})=\widetilde{M}\sslash_{0}SO(3)=\mu^{-1}(0)/SO(3)$. Notice that there is a Hamiltonian $S^{1}$-action on $\widetilde{M}$ with moment map $\widetilde{\ell_{12}}(\rho)=|\rho_{1}+\rho_{2}|$, and this $S^{1}$-action commutes with the $SO(3)$-action on $\widetilde{M}$. Hence, we can identify $\ell_{12}^{-1}(c)/S^{1}$ with $F^{-1}(c,0)/(S^{1}\times SO(3))$, where $F=(\widetilde{\ell_{12}},\mu):\widetilde{M}\to{\mathbb{R}}\times{\mathbb{R}}^{3}$. It follows that we can do symplectic reductions by stages: first by the $S^{1}$-action and then by the $SO(3)$-action. In other words, $\ell_{12}^{-1}(c)/S^{1}=(\widetilde{M}\sslash_{c}S^{1})\sslash_{0}SO(3)$.
By Lemma 2.1, $\widetilde{\ell_{12}}^{-1}(c)/S^{1}\cong S^{2}_{c}\times S^{2}_{r_{3}}\times\cdots\times S^{2}_{r_{n}}$. Hence, by definition of polygon spaces, $\ell_{12}^{-1}(c)/S^{1}=(\widetilde{M}\sslash_{c}S^{1})\sslash_{0}SO(3)=(S^{2}_{c}\times S^{2}_{r_{3}}\times\cdots\times S^{2}_{r_{n}})\sslash_{0}SO(3)=\mathcal{P}(c,r_{3},\ldots,r_{n}).$
[…]
Lemma 2.3.
Let $I,J$ be nonempty subsets of $\{1,\ldots,n\}$. If one of the following holds
  • $I,J$ are disjoint, or
  • $I^{c},J^{c}$ are disjoint, or
  • one of $I,J$ is contained in another,
then $\{\ell_{I}^{2},\ell_{J}^{2}\}=0$. Moreover, if $\ell_{I},\ell_{J}$ never vanish on $\mathcal{P}(r_{1},\ldots,r_{n})$, then $\{\ell_{I},\ell_{J}\}=0$.
We next establish a general criterion characterizing the singular points of $H_{t}$ that will be used in our semitoric families.
Lemma 2.4.
Fix $p\in\mathcal{P}(r_{1},\ldots,r_{n})$ and choose a representative $\rho\in\widetilde{\mathcal{P}}(r_{1},\ldots,r_{n})$ of $p$ with $\rho_{1}=(r_{1},0,0)$. For any $t\in[0,1]$, $p$ is a singular point of $H_{t}:=t\ell_{n-2,n-1}^{2}+(1-t)\ell^{2}_{n-1,n}$ if and only if there exists an $a\in{\mathbb{R}}$ such that
  • $(a,0,0)\times\rho_{i}=0$ for $i\neq n-2,n-1,n$,
  • $(a,0,0)\times\rho_{n-2}=t\rho_{n-1}\times\rho_{n-2}$, and
  • $(a,0,0)\times\rho_{n}=(1-t)\rho_{n-1}\times\rho_{n}$.
Proof.
By [KM96, Lemma 3.5], the Hamiltonian vector field $X_{H_{t}}$ of $H_{t}$ on $\widetilde{\mathcal{P}}(r_{1},\ldots,r_{n})$ is given by
$X_{H_{t}}(\rho)=(0,\ldots,0,t\rho_{n-1}\times\rho_{n-2},t\rho_{n-2}\times\rho_{n-1}+(1-t)\rho_{n}\times\rho_{n-1},(1-t)\rho_{n-1}\times\rho_{n}).$
Since the tangent space $T_{p}\mathcal{P}(r_{1},\ldots,r_{n})$ is the quotient of $T_{\rho}\widetilde{\mathcal{P}}(r_{1},\ldots,r_{n})$ by the vertical space $V_{\rho}$ generated by the infinitesimal vector fields of the $SO(3)$-action, $p$ is a singular point of $H_{t}$ if and only if $X_{H_{t}}(\rho)\in V_{\rho}$. It is straightforward to compute that
$V_{\rho}=\{((x,y,z)\times\rho_{1},\ldots,(x,y,z)\times\rho_{n}):x,y,z\in{\mathbb{R}}\}.$
Hence, $X_{H_{t}}\in V_{\rho}$ if and only if there exist $a,b,c\in{\mathbb{R}}$ such that $(a,b,c)\times\rho_{i}=0\textrm{ for }1\leq i\leq n-3$, $(a,b,c)\times\rho_{n-2}=t\rho_{n-1}\times\rho_{n-2}$ and $(a,b,c)\times\rho_{n}=(1-t)\rho_{n-1}\times\rho_{n}$. Since $\rho_{1}=(r_{1},0,0)$, this implies that $b=c=0$. The result follows immediately. ∎
Lemma 2.5.
Let $r_{1},\ldots,r_{5}$ be positive numbers such that
  • $r_{4}<r_{5}<r_{3}$
  • $|r_{1}+r_{2}-r_{3}|<r_{5}-r_{4}$
  • $|r_{3}-r_{4}-r_{5}|<|r_{1}-r_{2}|<r_{3}+r_{4}-r_{5}$
Then the $S^{1}$-action on $M:=\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ with moment map $J=\ell_{12}$ is semi-free such that $M^{S^{1}}$ has three connected components: two fixed surfaces $J^{-1}(|r_{1}-r_{2}|),J^{-1}(r_{1}+r_{2})$ and an isolated fixed point $p\in M$ with $J(p)=r_{3}+r_{4}-r_{5}$. Moreover, $p$ is also a singular point of $H_{t}:=t\ell_{34}^{2}+(1-t)\ell_{45}^{2}$ for any $t\in[0,1]$.
Proof.
By Lemma 2.3 and [KM96, Proposition 3.11] , $(\ell_{12},\ell_{34})$ (or $(\ell_{12},\ell_{45})$) is the moment map of an effective Hamiltonian $T^{2}$-action on $M$. It is straightforward to check that the Delzant polygons corresponding to this toric action are given by Figure 1. The claims about fixed components follow directly from Delzant’s classification [Del88] and the local normal form theorem [GS84]. Let $p\in M$ be the preimage of $(r_{3}+r_{4}-r_{5},r_{3}+r_{4})$ and let $\rho\in\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ be a representative of $p$. Since $\ell_{34}(p)=r_{3}+r_{4}$, we know that $\rho_{3}\times\rho_{4}=0$. Since $\rho_{1}+\rho_{2}=-\rho_{3}-\rho_{4}-\rho_{5}$ and $\ell_{12}(p)=r_{3}+r_{4}-r_{5}$, we know that $\rho_{4}\times\rho_{5}=0$ as well. Hence, by Lemma 2.4, $p$ is a singular point of $H_{t}$ for all $t\in[0,1]$.
$(J_{\min},r_{3}-r_{4})$$(J_{\max},r_{3}-r_{4})$$(J_{\max},r_{3}+r_{4})$$(j,r_{3}+r_{4})$$(J_{\min},r_{5}+J_{\min})$
(a) moment image of $(\ell_{12},\ell_{34})$
$(J_{\min},r_{4}+r_{5})$$(J_{\max},r_{4}+r_{5})$$(J_{\max},r_{5}-r_{4})$$(j,r_{5}-r_{4})$$(J_{\min},r_{3}-J_{\min})$
(b) moment image of $(\ell_{12},\ell_{45})$
Figure 1. In both figures, $J_{\min}=|r_{1}-r_{2}|,J_{\max}=r_{1}+r_{2},j=r_{3}+r_{4}-r_{5}$.
3. Semitoric Systems
In this section, we recall the definition of semitoric systems and the classification of non-degenerate singularities. The non-degeneracy of singularities can be defined in any dimension for any integrable system. Since our main interest lies in dimension 4, we will follow Section 2.3 in [LFP23] to give the equivalent definition using Hessian matrices. For the definition in any dimension using Cartan subalgebras, we refer the readers to [BF04, Definition 1.23].
3.1. Singularities of $4$-dimensional integrable systems
Let $(M,\omega)$ be a $4$-dimensional symplectic manifold. Recall that an integrable system on $(M,\omega)$ is a pair of functions $F=(J,H):M\to{\mathbb{R}}^{2}$ such that $dJ,dH$ are linearly independent almost everywhere and $\{J,H\}=0$. We say a point $p\in M$ is singular if $dJ_{m}$ and $dH_{m}$ are linearly dependent; the rank of the singular point $p$ is the dimension of the vector space spanned by $dJ_{m},dH_{m}$.
Rank zero points
Let $p\in M$ be a rank zero singular point and fix a basis of $T_{p}M$. Let $\Omega_{p}$ be the matrix of the symplectic form $\omega_{p}$ and $\operatorname{Hess}(J)_{p},\operatorname{Hess}(H)_{p}$ be the Hessian matrices under this basis. For any $\nu,\mu\in{\mathbb{R}}$, the characteristic polynomial of the matrix $A_{\nu,\mu}:=\Omega_{p}^{-1}(\nu\operatorname{Hess}(J)_{p}+\mu\operatorname{Hess}(H)_{p})$ is of the form $X\mapsto\chi_{\nu,\mu}(X^{2})$ for some quadratic polynomial $\chi_{\nu,\mu}$ (c.f. Proposition 1.2 and the discussion below Theorem 1.3 in [BF04]).
Definition 3.1.
Let $p\in M$ be a rank zero singular point of the integrable system $(J,H)$ on a 4-dimensional symplectic manifold $(M,\omega)$. $p$ is non-degenerate if and only if there exist $\nu,\mu\in{\mathbb{R}}$ such that $\chi_{\nu,\mu}$ has two distinct nonzero roots $\lambda_{1},\lambda_{2}\in{\mathbb{C}}$. In this case, the type of $p$ is
  • elliptic-elliptic if $\lambda_{1},\lambda_{2}<0$;
  • elliptic-hyperbolic if $\lambda_{1}<0,\lambda_{2}>0$;
  • hyperbolic-hyperbolic if $\lambda_{1},\lambda_{2}>0$;
  • focus-focus if both $\lambda_{1},\lambda_{2}$ are not purely real numbers, i.e. $\mathcal{I}(\lambda_{1})\neq 0$ and $\mathcal{I}(\lambda_{2})\neq 0$.
Rank one points
Let $p\in M$ be a rank one singular point. In this case, $dJ_{p}$ and $dH_{p}$ are linearly dependent, so the Hamiltonian vector fields $X_{J}(p),X_{H}(p)$ are linearly dependent as well because of the non-degeneracy of $\omega$. Let $L$ be the real span of $X_{J}(p),X_{H}(p)$ and $L^{\omega}$ be the symplectic orthogonal of $L$. Then $A_{\nu,\mu}:=\Omega_{p}^{-1}(\nu\operatorname{Hess}(J)_{p}+\mu\operatorname{Hess}(H)_{p})$ descends to an operator, by a slight abuse of notation, still called $A_{\nu,\mu}$ on the two-dimensional quotient $L^{\omega}/L$, and the eigenvalues of $A_{\nu,\mu}$ are of the form $\pm\lambda$ for $\lambda\in{\mathbb{C}}$ (again, see Proposition 1.2 in [BF04]).
Definition 3.2.
Let $p\in M$ be a rank one singular point of the integrable system $(J,H)$ on a 4-dimensional symplectic manifold $(M,\omega)$. $p$ is non-degenerate if and only if $A_{\nu,\mu}$ is an isomorphism on the quotient space $L^{\omega}/L$. In this case, the type of $p$ is
  • elliptic-regular if $A_{\nu,\mu}$ has eigenvalues $\pm i\alpha$ for some $\alpha\in{\mathbb{R}}\setminus\{0\}$;
  • hyperbolic-regular if $A_{\nu,\mu}$ has eigenvalues $\pm\alpha$ for some $\alpha\in{\mathbb{R}}\setminus\{0\}$.
Now, we are ready to define semitoric systems.
Definition 3.3.
A semitoric system is a 4-dimensional integrable system $(M,\omega,F=(J,H))$ such that
  1. (1)
    $J$ is proper,
  2. (2)
    $J$ is the moment map for an effective Hamiltonian $S^{1}$-action, and
  3. (3)
    all singularities of $F$ are non-degenerate without hyperbolic types.
Notice that (3) is equivalent to saying that all rank zero singular points are of either elliptic-elliptic or focus-focus type, and all rank one singular points are of elliptic-regular type.
3.2. Local normal forms of singularities of a semitoric system
In this section, we recall the local theory of singularities of a semitoric system. The best-known local theory is proposed by Eliasson [Eli84]: it is a symplectic Morse lemma for any integrable system. However, since we are studying semitoric systems only, the local normal form given by the Hamiltonian $S^{1}$-action comes in handy, so we will use the following lemma. We refer the interested readers to Section 2.2 in [LFP23] for a discussion of Eliasson’s normal form and the references therein.
3.2.1. Rank zero points
We first discuss the local forms around rank zero points. Notice that these points are all fixed by the $S^{1}$-action.
Lemma 3.4.
[LFP23, Proposition 7.5] Let $J=\epsilon(\frac{1}{2}|z_{1}|^{2}-\frac{1}{2}|z_{2}|^{2})$ with $\epsilon\in\{-1,1\}$. Then $H\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ is a Hamiltonian such that $\{J,H\}=0$, $H(0)=0$ and $dH(0)=0$ if and only if there exist $\mu_{1},\mu_{2},\mu_{3},\psi\in{\mathbb{R}}$ and $R_{3}\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ with $R_{3}(z_{1},z_{2})=O(3)$ and $\{J,R_{3}\}=0$ such that
$H(z_{1},z_{2})=\mu_{1}\mathfrak{R}(e^{i\psi}z_{1}z_{2})+\mu_{2}|z_{1}|^{2}+\mu_{3}|z_{2}|^{2}+R_{3}(z_{1},z_{2}).$
Moreover, if $H$ is of this form and if $(\mu_{1},\mu_{2}+\mu_{3})\neq(0,0)$, then $(J,H)$ is integrable. In this case, the singular point $(0,0)$ of $(q_{1},H)$ is
  • of focus-focus type if $|\mu_{2}+\mu_{3}|<|\mu_{1}|$;
  • of elliptic-elliptic type if $|\mu_{2}+\mu_{3}|>|\mu_{1}|$;
  • degenerate if $|\mu_{2}+\mu_{3}|=|\mu_{1}|$.
In this statement $O(3)$ means $O(||z_{1},z_{2}||^{3})$; we will use this notation throughout the paper.
For a point $p\in M$ fixed by the $S^{1}$-action, the isotropy weights are the weight vectors of the isotropy representation of $S^{1}$ on $T_{p}M$. In dimension $4$, there are two (counted with multiplicity) isotropy weights at each fixed point. The above lemma discusses the case where the isotropy weights are $\pm 1$. In fact, this is the only case where $p$ is possibly a focus-focus singularity of the semitoric system. Another case of interest is where $p$ lies on a surface fixed by the $S^{1}$-action. Since the action is effective, the isotropy weights must be $0,1$ or 0,10,-1. We prove the following lemma assuming that $p$ has isotropy weights $0,1$. Notice that we could have used the complex coordinates $z_{1},z_{2}$ and put $J$ as $|z_{1}|^{2}$, but this will put $H$ in a complicated form. Since the non-degeneracy and the type of singularity only depend on the second-order terms in the Taylor expansion, we use the real coordinates to simplify the statement and the proof.
Lemma 3.5.
Consider the symplectic vector space $({\mathbb{R}}^{4},\omega=dx_{1}\wedge dy_{1}+dx_{2}\wedge dy_{2})$. Let $J(x_{1},y_{1},x_{2},y_{2})=x_{1}^{2}+y_{1}^{2}$ and $H\in\mathcal{C}^{\infty}({\mathbb{R}}^{4},{\mathbb{R}})$ be a smooth function. Then $\{J,H\}=0$, $H(0)=0$ and $dH(0)=0$ if and only if there exist $a,b,c,d\in{\mathbb{R}}$, $R_{3}\in\mathcal{C}^{\infty}({\mathbb{R}}^{4},{\mathbb{R}})$ with $R_{3}=O(3)$ such that $H=a(x_{1}^{2}+y_{1}^{2})+bx_{2}^{2}+cx_{2}y_{2}+dy_{2}^{2}+R_{3}(x_{1},y_{1},x_{2},y_{2})$ and $\{J,R_{3}\}=0$. Moreover, in this case, $(0,0)$ is an elliptic-elliptic (elliptic-hyperbolic) point of $(J,H)$ if and only if $(0,0)$ is an elliptic (hyperbolic) singular point of $H|_{J=0}$ in the Morse sense.
Proof.
It is clear that if $H$ is in the given form, $\{J,H\}=0$, $H(0)=0$ and $dH(0)=0$. Conversely, we take the Taylor expansion of $H$. Since $H(0)=0$ and $dH(0)=0$, $H$ has no constant or linear terms. Hence, we can write $H$ as
$H=\sum_{i\leq j}a_{ij}x_{i}x_{j}+\sum_{i,j}b_{ij}x_{i}y_{j}+\sum_{i\leq j}c_{ij}y_{i}y_{j}+R_{3}(x_{1},y_{1},x_{2},y_{2}).$
It is straightforward to check that
$\{J,H\}=2b_{11}x_{1}^{2}+2b_{21}x_{1}x_{2}+4(c_{11}-a_{11})x_{1}y_{1}+2c_{12}x_{1}y_{2}-2a_{12}x_{2}y_{1}-2b_{11}y_{11}^{2}-2b_{12}y_{1}y_{2}+\{J,R_{3}\}.$
Since $\{J,R_{3}\}$ contains terms of order at least $3$, $\{J,H\}=0$ implies that
$a_{11}=c_{11},a_{12}=b_{11}=b_{12}=b_{21}=c_{12}=0,\textrm{ and }\{J,R_{3}\}=0.$
Therefore, $H=a_{11}(x_{1}^{2}+y_{1}^{2})+a_{22}x_{2}^{2}+b_{22}x_{2}y_{2}+c_{22}y_{2}^{2}+R_{3}(x_{1},y_{1},x_{2},y_{2})$ and $\{J,R_{3}\}=0$.
In this case, we have the following matrices
$\Omega=\begin{pmatrix}\mathbf{J}&\mathbf{O}\\ \mathbf{O}&\mathbf{J}\end{pmatrix},\operatorname{Hess}(J)=\begin{pmatrix}2\mathbf{I}&\mathbf{O}\\ \mathbf{O}&\mathbf{O}\end{pmatrix},\operatorname{Hess}(H)=\begin{pmatrix}2a_{11}\mathbf{I}&\mathbf{O}\\ \mathbf{O}&\mathbf{D}\end{pmatrix}$,
where $\mathbf{J}=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$, $\mathbf{D}=\begin{pmatrix}2a_{22}&b_{22}\\ b_{22}&2c_{22}\end{pmatrix}$, $\mathbf{O}$ is the zero matrix and $\mathbf{I}$ is the identity matrix. Fix a number $\lambda\neq-a_{11}$ and consider $A_{\lambda,1}=\Omega^{-1}(\lambda\operatorname{Hess}(J)+\operatorname{Hess}(H))$. The characteristic polynomial of $A_{\lambda,1}$ is $x^{2}+(2\lambda+2a_{11})^{2}][x^{2}+4a_{22}c_{22}-{b_{22}}^{2}]$. Hence, $\chi_{\lambda,1}(X)=(X+(2\lambda+2a_{11})^{2})(X+4a_{22}c_{22}-{b_{22}}^{2})$. The two roots are two real numbers $\lambda_{1}=-(2\lambda+2a_{11})^{2}$ and $\lambda_{2}={b_{22}}^{2}-4a_{22}c_{22}$. Since $\lambda\neq-a_{11}$, we have that $\lambda_{1}<0$. By Definition 3.1, $(0,0)$ is elliptic-elliptic if and only if $\lambda_{2}<0$, and elliptic-hyperbolic if and only if $\lambda_{2}>0$. Finally, we notice that $\mathbf{D}$ is the Hessian matrix of $H|_{J=0}$, so $(0,0)$ is an elliptic (hyperbolic) singular point in the Morse sense if and only if $\det(\mathbf{D})=-\lambda_{2}>0$ ($\det(\mathbf{D})=-\lambda_{2}<0$). The claim follows immediately. ∎
Recall that in a semitoric transition family, the type of singularity of the transition point changes as the parameter varies. The following proposition gives a criterion for when a transition point exists and proves that the transition times are roots of a quadratic polynomial.
Proposition 3.6.
Let $J=\frac{1}{2}|z_{1}|^{2}-\frac{1}{2}|z_{2}|^{2}$. Fix $\mu_{1},\mu_{2},\mu_{3},\psi,\nu_{2},\nu_{3}\in{\mathbb{R}}$ and $Q_{3},R_{3}\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ with $Q_{3}(z_{1},z_{2})=O(3),R_{3}(z_{1},z_{2})=O(3)$ and $\{J,Q_{3}\}=\{J,R_{3}\}=0$. Define $H_{0}(z_{1},z_{2}):=\nu_{2}|z_{1}|^{2}+\nu_{3}|z_{2}|^{2}+Q_{3}(z_{1},z_{2})$ and $H_{1}(z_{1},z_{2}):=\mu_{1}\mathfrak{R}(e^{i\psi}z_{1}z_{2})+\mu_{2}|z_{1}|^{2}+\mu_{3}|z_{2}|^{2}+R_{3}(z_{1},z_{2})$. Suppose the following hold:
  • $H_{0}(0,0)$ is a non-degenerate local minimum in the Morse-Bott sense;
  • $H_{1}(0,0)$ is a non-degenerate local maximum in the Morse-Bott sense;
  • $(\mu_{2}+\mu_{3})^{2}>\mu_{1}^{2}>0$.
Then for $t\in[0,1]$, $(J,H_{t}:=(1-t)H_{0}+tH_{1})$ is integrable, and there exist $0<t^{-}<t^{+}<1$ such that the singular point $(0,0)$ of $(J,H_{t})$ is
  • of elliptic-elliptic type when $t\in[0,t^{-})\cup(t^{+},1]$,
  • of focus-focus type when $t\in(t^{-},t^{+})$, and
  • degenerate when $t=t^{-},t^{+}$.
Proof.
Since $H_{0}$ has a non-degenerate local minimum (in the Morse-Bott sense) at the origin, the quadratic form $\nu_{2}|z_{1}|^{2}+\nu_{3}|z_{2}|^{2}$ must be positive semi-definite and nonzero. This implies that $\nu_{2}+\nu_{3}>0$. Moreover, since $H_{1}$ has a non-degenerate local maximum (in the Morse-Bott sense) at the origin, $\operatorname{Hess}(H_{1})$ has non-positive eigenvalues with at least one strictly negative eigenvalues, so the sum of eigenvalues of $\operatorname{Hess}(H_{1})$ is strictly negative. Since $\text{det}\big(\lambda I-\text{Hess}(H_{1})\big)=\big(\lambda^{2}-2(\mu_{2}+\mu_{3})\lambda+4\mu_{2}\mu_{3}-\mu_{1}^{2}\big)^{2}$, it follows that $\mu_{2}+\mu_{3}<0$.
Consider the function $H_{t}=(1-t)H_{0}+tH_{1}=t\mu_{1}\mathfrak{R}(e^{i\psi}z_{1}z_{2})+\big(t\mu_{2}+(1-t)\nu_{2}\big)|z_{1}|^{2}+\big(t\mu_{3}+(1-t)\nu_{3}\big)|z_{2}|^{2}+O(3)$. When $t\neq 0$, we have $t\mu_{1}\neq 0$, so by Lemma 3.4 $(J,H_{t})$ is integrable for $t\in(0,1]$. Since $\nu_{2}+\nu_{3}>0$, $(J,H_{0})$ is integrable by the same lemma. Moreover, by Lemma 3.4 the singularity type of $(J,H_{t})$ at $(0,0)$ is determined entirely by the sign of
$f(t):=\big(t\mu_{2}+(1-t)\nu_{2}+t\mu_{3}+(1-t)\nu_{3}\big)^{2}-(t\mu_{1})^{2}.$
Indeed, the point is elliptic–elliptic whenever $f(t)>0$, focus–focus whenever $f(t)<0$, and degenerate whenever $f(t)=0$. Thus, the problem is reduced to analyzing the function $f(t)$ on the interval $[0,1]$.
So far, we have the following inequalities:
  1. (1)
    $(\mu_{2}+\mu_{3})^{2}>\mu_{1}^{2}$,
  2. (2)
    $\mu_{1}^{2}>0$,
  3. (3)
    $\nu_{2}+\nu_{3}>0$, and
  4. (4)
    $\mu_{2}+\mu_{3}<0$.
We can rewrite $f(t)$ as the following quadratic function in $t$:
$f(t)=\{\big((\mu_{2}+\mu_{3})-(\nu_{2}+\nu_{3})\big)^{2}-\mu_{1}^{2}\}t^{2}+2\big((\mu_{2}+\mu_{3})-(\nu_{2}+\nu_{3})\big)(\nu_{2}+\nu_{3})t+(\nu_{2}+\nu_{3})^{2}.$
By (1),(3), and (4), the coefficient of $t^{2}$ in $f(t)$, which is $\big((\mu_{2}+\mu_{3})-(\nu_{2}+\nu_{3})\big)^{2}-\mu_{1}^{2}=\{(\mu_{2}+\mu_{3})^{2}-\mu_{1}^{2}\}-2(\mu_{2}+\mu_{3})(\nu_{2}+\nu_{3})+(\nu_{2}+\nu_{3})^{2}$, is strictly positive. By (2) and (3), its discriminant $\Delta=4\mu_{1}^{2}(\nu_{2}+\nu_{3})^{2}$ is strictly positive. Hence $f(t)$ is a concave up parabola with two distinct real roots. Finally, we check their position relative to the interval $[0,1]$. At the left endpoint, by (3) and (4), we have
$f(0)=(\nu_{2}+\nu_{3})^{2}>0\quad\text{and}\quad f^{\prime}(0)=2\big((\mu_{2}+\mu_{3})-(\nu_{2}+\nu_{3})\big)(\nu_{2}+\nu_{3})<0,$
At the right endpoint, by (1), (3) and (4), we have
$f(1)=(\mu_{2}+\mu_{3})^{2}-\mu_{1}^{2}>0\quad\text{and}\quad f^{\prime}(1)=2[\{(\mu_{2}+\mu_{3})^{2}-\mu_{1}^{2}\}-(\mu_{2}+\mu_{3})(\nu_{2}+\nu_{3})]>0.$
Hence, there exist $t^{-},t^{+}$ such that $0<t^{-}<t^{+}<1$ with $f(t^{\pm})=0$, and we have
[…]
By Lemma 3.4, $(0,0)$ is a singular point of the elliptic–elliptic type for $t\in[0,t^{-})\cup(t^{+},1]$, focus-focus type for $t\in(t^{-},t^{+})$, and degenerate at $t^{\pm}$. This shows the claimed transition between elliptic-elliptic and focus-focus singularity. ∎
3.2.2. Rank one points
Recall that a point $p\in M$ has rank one if $dJ_{p},dH_{p}$ span a one-dimensional vector space. When $dJ_{p}\neq 0$, $p$ is not a fixed point of the $S^{1}$-action. If $S^{1}$ acts freely on $p$, we can apply symplectic reduction (at least locally) at the level $J(p)$ and study the type of singularity of $p]\in M^{\textrm{red},J(p)}=J^{-1}(J(p))/S^{1}$ of the function $H^{\textrm{red},J(p)}$ to which $H$ descends.
Lemma 3.7.
[LFP23, Lemma 2.6] Let $(M,\omega,F=(J,H))$ be a 4-dimensional integrable system, where $J$ is the moment map of an effective Hamiltonian $S^{1}$-action on $M$. Let $p\in M$ be a point on which $S^{1}$ acts freely. Let $j=J(p)$. Then $p$ is a rank one singular point of $F$ if and only if $[p]$ is a singular point of $H^{\textrm{red},j}$. Moreover, $p$ is a non-degenerate singular point of $F$ if and only if $[p]$ is non-degenerate for $H^{\textrm{red},j}$ (in the Morse sense) and the type of $p$ is elliptic-regular (respectively hyperbolic-regular) if and only if $[p]$ is an elliptic (respectively hyperbolic) singular point of $H^{\textrm{red},j}$.
When $dJ_{p}=0$ and $dH_{p}\neq 0$, $p$ lies on a fixed surface. In this case, $p$ is automatically an elliptic-regular singularity.
Lemma 3.8.
Consider the symplectic vector space $({\mathbb{C}}^{2},\omega=\frac{i}{2}(dz_{1}\wedge d\bar{z_{1}}+dz_{2}\wedge d\bar{z_{2}}))$. Let $J(z_{1},z_{2})=|z_{1}|^{2}$ and $H\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ be a smooth function. Then $\{J,H\}=0$, $H(0)=0$ and $dH(0)\neq 0$ if and only if there exist $a,b\in{\mathbb{R}}$ not both zero, $R_{2}\in\mathcal{C}^{\infty}({\mathbb{C}}^{2},{\mathbb{R}})$ with $R_{2}=O(2)$ such that $H=a\mathcal{R}(z_{2})+b\mathcal{I}(z_{2})+R_{2}(z_{1},z_{2})$ and $\{J,R_{2}\}=0$. Moreover, in this case, $(0,0)$ is an elliptic-regular point of $(J,H)$.
Proof.
The proof of the first claim is a similar computation to the proof of Lemma 3.5, so we leave it to the readers to verify. Since $X_{J}(0)=0$, the real span of $X_{J}(0),X_{H}(0)$ is the same as the real span of $X_{H}(0)$. Let $L$ denote this real span. Then $L^{\omega}={\mathbb{R}}(\frac{\partial}{\partial z_{1}},\frac{\partial}{\partial\overline{z}_{1}},X_{H}(0))$. Hence, we can use the equivalent classes of $\frac{\partial}{\partial z_{1}},\frac{\partial}{\partial\overline{z}_{1}}$ as a basis of $L^{\omega}/L$. Under this basis, we have the following matrix representation:
$\Omega=\begin{pmatrix}0&\frac{i}{2}\\ -\frac{i}{2}&0\end{pmatrix}\hskip 56.9055pt\operatorname{Hess}(J)=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$
Hence, $A_{1,0}=\Omega^{-1}\operatorname{Hess}(J)$ has eigenvalues $\pm 2i$, so by definition, $(0,0)$ is an elliptic-regular point of $(J,H)$. ∎
Lemma 3.5 and Lemma 3.8 together imply the following non-degeneracy result for points on a fixed surface.
Proposition 3.9.
Let $(M,\omega,(J,H))$ be a 4-dimensional integrable system, where $J$ is the moment map of an effective Hamiltonian $S^{1}$-action on $M$. Let $\Sigma\subset M^{S^{1}}$ be a fixed surface. Then any rank one singular point $p\in\Sigma$ is of elliptic-regular type. Moreover, a rank zero singular point $p\in\Sigma$ is of elliptic-elliptic type (elliptic-hyperbolic type respectively) if and only if $p$ is an elliptic (hyperbolic resp.) singular point of $H|_{\Sigma}$ in the Morse sense.
4. Non-degeneracy of non-transition points
We will prove our main result in the remaining sections. Let $P\in\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ be defined as in the statement of Theorem 1.1. In this section, we check the non-degeneracy of singular points in $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})\setminus\{P\}$. We start by proving a helpful lemma.
Lemma 4.1.
Fix positive numbers $c,r_{3},r_{4},r_{5}$ such that $\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a smooth manifold and $r_{3}>r_{5}$. For $t\in(0,1)$, define $H_{t}:\mathcal{P}(c,r_{3},r_{4},r_{5})\to{\mathbb{R}}$ by $H_{t}([(\rho_{0},\rho_{3},\rho_{4},\rho_{5})])=(1-t)|\rho_{4}+\rho_{5}|^{2}+t|\rho_{3}+\rho_{4}|^{2}$. Then $Q\in\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a singular point of $H_{t}$ if and only if there exists a representative $\rho$ with $\rho_{0}=(c,0,0),\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$ and a nonzero number $a\in{\mathbb{R}}$ such that
(*) $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})\textrm{ and }ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4}).$
In this case, the following holds:
  1. (1)
    $b_{3},b_{5}$ are nonzero. When $t\neq\dfrac{1}{2}$, $b_{4}$ is also nonzero.
  2. (2)
    When $t\neq\dfrac{1}{2}$ or $b_{4}\neq 0$, we have $K:=\dfrac{tb_{5}}{(1-t)b_{3}}\neq-1$. When $t=\dfrac{1}{2}$ and $b_{4}=0$, we have $a_{3}b_{5}-a_{5}b_{3}\neq 0$.
Proof.
Consider a representative $\rho=(\rho_{0},\rho_{3},\rho_{4},\rho_{5})\in\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ of $Q$ such that $\rho_{0}=(c,0,0)$. By Lemma 2.4, $Q$ is a singular point if and only if there exists an $a\in{\mathbb{R}}$ such that $(a,0,0)\times\rho_{3}=t\rho_{4}\times\rho_{3}$ and $(a,0,0)\times\rho_{5}=(1-t)\rho_{4}\times\rho_{5}$. This implies that $\rho_{0},\rho_{3},\rho_{4},\rho_{5}$ are coplanar, so we can further assume that $\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$. If $a=0$, then we have that $\rho_{3},\rho_{4},\rho_{5}$ are colinear. Since $\rho_{0}+\rho_{3}+\rho_{4}+\rho_{5}=0$, this implies that $\rho_{0},\rho_{3},\rho_{4},\rho_{5}$ are all colinear, which is impossible since $\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a smooth manifold. Hence, we conclude that $Q$ is a singular point if and only if there exists a representative $\rho$ of the form $\rho_{0}=(c,0,0),\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$ such that there exists an $a\in{\mathbb{R}}\setminus\{0\}$ with
$ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})\textrm{ and }ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4}).$
(1) Regardless of the value of $t$, if $b_{3}=0$, then the equation $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})$ implies that $b_{4}=0$. However, since $\rho_{0}+\rho_{3}+\rho_{4}+\rho_{5}=0$, this further implies that $b_{5}=0$ and thus $\rho_{0},\rho_{3},\rho_{4},\rho_{5}$ are all colinear, which is impossible, so $b_{3}\neq 0$. Similarly, $b_{5}\neq 0$. Assume that $t\neq\dfrac{1}{2}$. If $b_{4}=0$, then Equation (*4.1) implies that $a=ta_{4},a=(1-t)a_{4}$, which is impossible unless $t=\dfrac{1}{2}$. Hence, we conclude that $b_{4}$ is nonzero whenever $t\neq\dfrac{1}{2}$.
(2) When $t\neq\dfrac{1}{2}$, we prove $K\neq-1$ by contradiction. If K=1K=-1, then $b_{5}=-\dfrac{1-t}{t}b_{3}$, so $b_{4}=-b_{3}-b_{5}=\dfrac{1-2t}{t}b_{3}=-\dfrac{1-2t}{1-t}b_{5}$. Since $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})$ and $b_{3}\neq 0$, we can substitute $b_{4}$ by $\dfrac{1-2t}{t}b_{3}$ and divide $b_{3}$ on both sides to get that
$a=t\bigg(a_{4}-a_{3}\frac{1-2t}{t}\bigg)=ta_{4}-(1-2t)a_{3}.$
Similarly, using $ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4})$ and $b_{4}=-\dfrac{1-2t}{1-t}b_{5}$, we obtain
$a=(1-t)\bigg(a_{4}+a_{5}\frac{1-2t}{1-t}\bigg)=(1-t)a_{4}+(1-2t)a_{5}.$
These two equations imply that $(1-2t)(a_{3}+a_{4}+a_{5})=0$. If $t\neq\dfrac{1}{2}$, $a_{3}+a_{4}+a_{5}=0$. This is impossible as $c+a_{3}+a_{4}+a_{5}=0$ and $c\neq 0$. Therefore, $K\neq-1$ whenever $t\neq\dfrac{1}{2}$.
When $t=\dfrac{1}{2}$, $K=\dfrac{b_{5}}{b_{3}}$. Since $b_{3}+b_{4}+b_{5}=0$, we have $1+\dfrac{b_{4}}{b_{3}}+K=0$. Hence, if $b_{4}\neq 0$, $K\neq-1$. If $b_{4}=0$, then $b_{3}=-b_{5}$, so $a_{3}b_{5}-a_{5}b_{3}=b_{5}(a_{3}+a_{5})$. Since $r_{3}>r_{5}$, we have $|a_{3}|=\sqrt{r_{3}^{2}-b_{3}^{2}}>\sqrt{r_{5}^{2}-b_{5}^{2}}=|a_{5}|$ and thus $a_{3}b_{5}-a_{5}b_{3}=b_{5}(a_{3}+a_{5})\neq 0$. ∎
The following proposition takes care of almost all singular points except for the transition point and singular points that have the same moment image as the transition point.
Proposition 4.2.
Let $r_{1},\ldots,r_{5}$ be positive numbers such that
  • $r_{4}<r_{5}<r_{3}$,
  • $|r_{1}+r_{2}-r_{3}|<r_{5}-r_{4}$,
  • $|r_{3}-r_{4}-r_{5}|<|r_{1}-r_{2}|<r_{3}+r_{4}-r_{5}$.
Let $M:=\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. Consider the family $(J,H_{t})=(\ell_{12},t\ell_{34}^{2}+(1-t)\ell_{45}^{2})$ on $M$, where $t\in(0,1)$. Fix $|r_{1}-r_{2}|\leq c\leq r_{1}+r_{2}$ such that $c\neq r_{3}+r_{4}-r_{5}$. Let $M^{\textrm{red},c}:=J^{-1}(c)/S^{1}$ be the symplectic reduction at level $c$ and let $H^{\textrm{red},c}_{t}$ be the function to which $H_{t}$ descends. Then $H^{\textrm{red},c}_{t}$ is a Morse function with only elliptic singularities.
Proof.
By Lemma 2.5, $M^{\textrm{red},c}$ is a smooth manifold if $c\neq|r_{1}-r_{2}|,r_{1}+r_{2},r_{3}+r_{4}-r_{5}$. Moreover, if $c=|r_{1}-r_{2}|$ or $c=r_{1}+r_{2}$, then $M^{\textrm{red},c}=J^{-1}(c)/S^{1}=J^{-1}(c)$ is a smooth surface fixed by the $S^{1}$-action. Hence, by Proposition 2.2, $M^{\textrm{red},c}\cong\mathcal{P}(c,r_{3},r_{4},r_{5})$.
Let $Q\in M^{\textrm{red},c}$ be a singular point of $H^{\textrm{red},c}_{t}$. By Lemma 4.1, there exists a representative $\rho$ of $Q$ with $\rho_{0}=(c,0,0),\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$ and a nonzero number $a\in{\mathbb{R}}$ such that
(*) $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})\textrm{ and }ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4}).$
To compute the Hessian at $Q$, we find $W_{1},W_{2}\in T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ such that $W_{1}],[W_{2}]$ form a basis of $T_{Q}M^{\textrm{red},c}$ and then find the matrix representation of the Hessian of $H_{t}^{\textrm{red},c}$ under this basis. Since $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ is a submanifold of $S^{2}_{c}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}$, it suffices to find a coordinate chart of $S^{2}_{c}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}$ around $\rho$. By item (1) of Lemma 4.1, we have $b_{3},b_{5}\neq 0$. Depending on whether $b_{4}$ is zero, we choose different charts.
Case 1: when $b_{4}\neq 0$, we choose the charts with coordinates $(y_{0},z_{0},x_{3},z_{3},x_{4},z_{4},x_{5},z_{5})$. In this chart, the infinitesimal vector field at $\rho$ generated by the $SO(3)$-action is
$V_{1}=b_{3}\frac{\partial}{\partial z_{3}}+b_{4}\frac{\partial}{\partial z_{4}}+b_{5}\frac{\partial}{\partial z_{5}},\ V_{2}=-c\frac{\partial}{\partial z_{0}}-a_{3}\frac{\partial}{\partial z_{3}}-a_{4}\frac{\partial}{\partial z_{4}}-a_{5}\frac{\partial}{\partial z_{5}},\ V_{3}=c\frac{\partial}{\partial y_{0}}-b_{3}\frac{\partial}{\partial x_{3}}-b_{4}\frac{\partial}{\partial x_{4}}-b_{5}\frac{\partial}{\partial x_{5}}.$
Let $V_{\rho}$ be the real span of $V_{1},V_{2},V_{3}$. Define $W_{1}=\frac{\partial}{\partial z_{3}}-\frac{\partial}{\partial z_{5}}$ and $W_{2}=(a_{5}b_{3}-a_{3}b_{5})\frac{\partial}{\partial y_{0}}-b_{3}b_{5}\frac{\partial}{\partial x_{3}}+b_{3}b_{5}\frac{\partial}{\partial x_{5}}$. It’s straightforward to verify that $W_{1},W_{2},V_{1},V_{2},V_{3}$ form a basis of $T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$. Since $T_{Q}M^{\textrm{red},c}\cong T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})/V_{\rho}$, $W_{1}],[W_{2}]$ form a basis of $T_{Q}M^{\textrm{red},c}$.
Recall that the Hessian of a function $f$ on a manifold $N$ at a singular point $n\in N$ can be computed in the following way: if $X$ and $Y$ are tangent vectors in $T_{n}N$, then $\operatorname{Hess}(f)_{n}(X,Y):=X\cdot(\widetilde{Y}\cdot f)(n)$, where $\widetilde{Y}$ denotes a vector field extending $Y$ locally.
Since $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ is a principal $SO(3)$-bundle over $\mathcal{P}(c,r_{3},r_{4},r_{5})$, $\operatorname{Hess}(H_{t}^{\textrm{red},c})_{Q}([W_{i}],[W_{j}])=\operatorname{Hess}(H_{t}^{\textrm{red},c})_{\rho}(W_{i},W_{j})$, and we will compute the latter by extending $W_{1},W_{2}$ locally on $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$. Let $\widetilde{W}_{i}$ be a local extension on $W_{i}$ for $i=1,2$. We can write it as
$\widetilde{W}_{i}=f^{i}\frac{\partial}{\partial y_{0}}+g^{i}\frac{\partial}{\partial z_{0}}+\sum_{j=3}^{5}f_{j}^{i}\frac{\partial}{\partial x_{j}}+\sum_{j=3}^{5}g_{j}^{i}\frac{\partial}{\partial z_{j}}$
Since $\widetilde{W}_{i}$ is tangent to $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$, the coefficient functions satisfy the following equations:
(1) $f^{i}\frac{y_{0}}{x_{0}}-g^{i}\frac{z_{0}}{x_{0}}+f_{3}^{i}+f_{4}^{i}+f_{5}^{i}=0$
(2) $f^{i}-\sum_{j=3}^{5}\frac{x_{j}}{y_{j}}f_{j}^{i}-\sum_{j=3}^{5}\frac{z_{j}}{y_{j}}g_{j}^{i}=0$
(3) $g^{i}+g_{3}^{i}+g_{4}^{i}+g_{5}^{i}=0$
It is straightforward to compute that
$W_{1}(\widetilde{W}_{i}H_{1}^{\textrm{red},c})(\rho)=\frac{a_{4}b_{3}-a_{3}b_{4}}{b_{3}}\left(\frac{\partial f_{3}^{i}}{\partial z_{3}}(\rho)-\frac{b_{3}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{3}}(\rho)-\frac{\partial f_{3}^{i}}{\partial z_{5}}(\rho)+\frac{b_{3}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{5}}(\rho)\right)-g_{3}^{i}(\rho)\frac{b_{4}}{b_{3}}$
$W_{1}(\widetilde{W}_{i}H_{0}^{\textrm{red},c})(\rho)=\frac{a_{4}b_{5}-a_{5}b_{4}}{b_{5}}\left(\frac{\partial f_{5}^{i}}{\partial z_{3}}(\rho)-\frac{b_{5}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{3}}(\rho)-\frac{\partial f_{5}^{i}}{\partial z_{5}}(\rho)+\frac{b_{5}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{5}}(\rho)\right)+g_{5}^{i}(\rho)\frac{b_{4}}{b_{5}}$
Since $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})$, $ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4})$ and $b_{3}+b_{4}+b_{5}=0$, we have
$W_{1}(\widetilde{W}_{i}H_{t}^{\textrm{red},c})(\rho)$ $=tW_{1}(\widetilde{W}_{i}H_{1}^{\textrm{red},c})(\rho)+(1-t)W_{1}(\widetilde{W}_{i}H_{0}^{\textrm{red},c})(\rho)$
$=a\left(\frac{\partial(f_{3}^{i}+f_{4}^{i}+f_{5}^{i})}{\partial z_{3}}(\rho)-\frac{\partial(f_{3}^{i}+f_{4}^{i}+f_{5}^{i})}{\partial z_{5}}(\rho)\right)-tg_{3}^{i}(\rho)\frac{b_{4}}{b_{3}}+(1-t)g_{5}^{i}(\rho)\frac{b_{4}}{b_{5}}$
By differentiating Equation (1) with respect to $z_{j}$ for $j=3,4,5$ and evaluating it at $\rho$, we have $\frac{\partial(f_{3}^{i}+f_{4}^{i}+f_{5}^{i})}{\partial z_{j}}(\rho)=0$. Hence, we have
$W_{1}(\widetilde{W}_{i}H_{t}^{\textrm{red},c})(\rho)=-tg_{3}^{i}(\rho)\frac{b_{4}}{b_{3}}+(1-t)g_{5}^{i}(\rho)\frac{b_{4}}{b_{5}},$
which evaluates to $\frac{b_{4}}{b_{3}}t-\frac{b_{4}}{b_{5}}(1-t)$ when $i=1$ and evaluates to 0 when $i=2$.
Similarly, one can compute
$W_{2}(\widetilde{W}_{2}H_{t}^{\textrm{red},c})(\rho)=\frac{a(a_{3}b_{5}-a_{5}b_{3})^{2}}{c}-t\frac{a_{3}^{2}+b_{3}^{2}}{b_{3}}b_{4}b_{5}^{2}-(1-t)\frac{a_{5}^{2}+b_{5}^{2}}{b_{5}}b_{4}b_{3}^{2}.$
Thus, the matrix representation of the Hessian of $H_{t}^{\textrm{red},c}$ at $Q$ under the basis $W_{1}],[W_{2}]$ is
$\begin{pmatrix}-t\frac{b_{4}}{b_{3}}-(1-t)\frac{b_{4}}{b_{5}}&0\\ 0&\frac{a}{c}(a_{3}b_{5}-a_{5}b_{3})^{2}-\frac{t}{b_{3}}(a_{3}^{2}+b_{3}^{2})b_{4}b_{5}^{2}-\frac{1-t}{b_{5}}(a_{5}^{2}+b_{5}^{2})b_{3}^{2}b_{4}\end{pmatrix}$
We show that this matrix has positive determinant. Set $X=a_{4}b_{3}-a_{3}b_{4},\ Y=a_{4}b_{5}-a_{5}b_{4}$, then $X+Y=a_{4}(b_{3}+b_{5})-(a_{3}+a_{5})b_{4}=-a_{4}b_{4}+(c+a_{4})b_{4}=cb_{4}$. Since $c\neq 0$, we can write $b_{4}=\frac{X+Y}{c}$. By Equation (*4.1), $ab_{3}=tX,ab_{5}=(1-t)Y$, so we have $X=\frac{ab_{3}}{t},Y=\frac{ab_{5}}{1-t}$. Hence,
$A:=-t\frac{b_{4}}{b_{3}}-(1-t)\frac{b_{4}}{b_{5}}=-b_{4}\left(\frac{t}{b_{3}}+\frac{1-t}{b_{5}}\right)=-\frac{X+Y}{c}\left(\frac{a}{X}+\frac{a}{Y}\right)=-\frac{a}{c}\left(2+\frac{X}{Y}+\frac{Y}{X}\right).$
Let $K=\frac{Y}{X}=\frac{tb_{5}}{(1-t)b_{3}}$, then $A=-\frac{a}{cK}(K+1)^{2}$. Similarly, one can compute
$B:=\frac{a}{c}(a_{3}b_{5}-a_{5}b_{3})^{2}-\frac{t}{b_{3}}(a_{3}^{2}+b_{3}^{2})b_{4}b_{5}^{2}-\frac{1-t}{b_{5}}(a_{5}^{2}+b_{5}^{2})b_{3}^{2}b_{4}=-\frac{a}{cK}\left[(Ka_{3}b_{5}+a_{5}b_{3})^{2}+(K+1)^{2}b_{3}^{2}b_{5}^{2}\right].$
By item (2) of Lemma 4.1, $K\neq-1$, so $A,B$ are both nonzero. Hence, $\det(\operatorname{Hess}(H_{t}^{\textrm{red},c}))=AB$ has the same sign as $\dfrac{a^{2}}{c^{2}K^{2}}>0$. We conclude that $Q$ is an elliptic singularity (in the Morse sense).
Case 2: when $b_{4}=0$, then $a_{4}\neq 0$, so we choose the charts with coordinates $(y_{0},z_{0},x_{3},z_{3},y_{4},z_{4},x_{5},z_{5})$. Notice that by (1) of Lemma 4.1, this could only happen when $t=\dfrac{1}{2}$. A similar computation shows that under the basis $W_{1}=\frac{\partial}{\partial z_{0}}-\frac{\partial}{\partial z_{3}}$ and $W_{2}=(a_{5}b_{3}-a_{3}b_{5})\frac{\partial}{\partial y_{0}}-b_{3}b_{5}\frac{\partial}{\partial x_{3}}+b_{3}b_{5}\frac{\partial}{\partial x_{5}}$, the matrix representation of the Hessian of $H_{t}^{\textrm{red},c}$ is given by
$\begin{pmatrix}\dfrac{a}{c}&0\\ 0&\dfrac{a}{c}(a_{3}b_{5}-a_{5}b_{3})^{2}\end{pmatrix}$
By (2) of Lemma 4.1, we know that $a_{3}b_{5}-a_{5}b_{3}\neq 0$, so the determinant of the matrix has the same sign as $\dfrac{a^{2}}{c^{2}}>0$. We conclude that $Q$ is an elliptic singularity (in the Morse sense). ∎
5. The transition point $P$
In this section, we study the transition point and the transition times.
Proposition 5.1.
Let $r_{1},\ldots,r_{5}$ be positive numbers such that
  • $\frac{1}{2}r_{5}\leq r_{4}<r_{5}<r_{3}$,
  • $|r_{1}+r_{2}-r_{3}|<r_{5}-r_{4}$, and
  • $|r_{3}-r_{4}-r_{5}|<|r_{1}-r_{2}|<r_{3}+r_{4}-r_{5}$.
Let $j=r_{3}+r_{4}-r_{5}$, and
$t^{\pm}=\frac{\,r_{3}r_{4}+r_{3}r_{5}+2r_{4}^{2}-3r_{4}r_{5}+r_{5}^{2}\,\pm\,2\sqrt{r_{3}r_{4}r_{5}j}}{\,(r_{3}+r_{5})^{2}+4r_{4}j\,}.$
Define $P\in\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ represented by the planar pentagon $\rho=(\rho_{1},\rho_{2},\rho_{3},\rho_{4},\rho_{5})$, where
$\rho_{1}=\left(\frac{r_{1}^{2}-r^{2}_{2}+j^{2}}{2j},\frac{\sqrt{\big((r_{1}+j)^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j)^{2}\big)}}{2j},0\right),$
$\rho_{2}=\left(\frac{r_{2}^{2}-r_{1}^{2}+j^{2}}{2j},-\frac{\sqrt{\big((r_{1}+j)^{2}-r_{2}^{2}\big)\big(r_{2}^{2}-(r_{1}-j)^{2}\big)}}{2j},0\right),$
$\rho_{3}=(-r_{3},0,0),\rho_{4}=(-r_{4},0,0),\rho_{5}=(r_{5},0,0).$
Then the following holds:
  • for $t\in[0,1]$, $P$ is a rank zero singular point of $(J,H_{t}):=(\ell_{12},t\ell_{34}^{2}+(1-t)\ell_{45}^{2})$,
  • for $t\in[0,t^{-})\cup(t^{+},1]$, $P$ is of elliptic-elliptic type, and
  • for $t\in(t^{-},t^{+})$, $P$ is of focus-focus type.
Proof.
We first notice that $\ell_{12}(P)=r_{3}+r_{4}-r_{5}$ and $\ell_{34}(P)=r_{3}+r_{4}$, so by Lemma 2.5, $P$ is a rank zero singular point of $(J,H_{t})$ for all $t\in[0,1]$. To check the type of singularity, we need to compute the Hessian matrix of $J,H_{t}$, and the matrix of the symplectic form under a basis. Since $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is a sub-quotient of $S^{2}_{r_{1}}\times S^{2}_{r_{2}}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}$, we can use the standard coordinates on $({\mathbb{R}}^{3})^{5}$ to induce the coordinates on $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. Since the $x$-coordinate of each edge of $\rho$ is nonzero, $(y_{1},z_{1},y_{2},z_{2},y_{3},z_{3},y_{4},z_{4},y_{5},z_{5})$ gives a coordinate chart on the product of spheres. Notice that $T_{P}\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})\cong T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})/V_{\rho}$, where $V_{\rho}$ is the space of infinitesimal vector fields generated by the $SO(3)$-action, so it suffices to find vectors in $T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ that descend to a basis of $T_{P}\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$.
Define $W_{1}=\frac{\partial}{\partial z_{1}}-\frac{\partial}{\partial z_{3}}$, $W_{2}=\frac{\partial}{\partial z_{1}}-\frac{\partial}{\partial z_{4}}$, $W_{3}=\frac{\partial}{\partial y_{3}}-\frac{\partial}{\partial y_{4}}$, $W_{4}=\frac{\partial}{\partial y_{4}}-\frac{\partial}{\partial y_{5}}$. Notice that $T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ is the kernel of $d\mu_{\rho}$, where $\mu:T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})\to{\mathbb{R}}^{3}$ is the moment map of the $SO(3)$-action, so $W_{i}\in T_{\rho}\widetilde{\mathcal{P}}(r_{1},r_{2},r_{3},r_{4},r_{5})$ for $i=1,2,3,4$. It is straightforward to verify that $W_{1},W_{2},W_{3},W_{4}$ span a four dimensional vector space that intersects $V_{\rho}$ trivially, so $W_{1}],[W_{2}],[W_{3}],[W_{4}]$ form a basis of $T_{P}\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. Using a similar calculation as in the proof of Proposition 4.2, the Hessian of $H_{0}$, $H_{1}$, and $J$ under this basis is given by
$\operatorname{Hess}(H_{0})_{P}=\begin{pmatrix}0&0&0&0\\ 0&\frac{r_{5}}{r_{4}}&0&0\\ 0&0&\frac{r_{5}}{r_{4}}&1-\frac{r_{5}}{r_{4}}\\ 0&0&1-\frac{r_{5}}{r_{4}}&\frac{r_{5}}{r_{4}}+\frac{r_{4}}{r_{5}}-2\end{pmatrix},\text{Hess}(H_{1})_{P}=\begin{pmatrix}-\frac{r_{4}}{r_{3}}&1&0&0\\ 1&-\frac{r_{4}}{r_{3}}&0&0\\ 0&0&-2-\frac{r_{4}}{r_{3}}-\frac{r_{3}}{r_{4}}&1+\frac{r_{3}}{r_{4}}\\ 0&0&1+\frac{r_{3}}{r_{4}}&-\frac{r_{3}}{r_{4}}\\ \end{pmatrix}$.
$\text{Hess}(J)_{P}=\begin{pmatrix}1-\frac{j}{r_{3}}&1&0&0\\ 1&1-\frac{j}{r_{4}}&0&0\\ 0&0&-j(\frac{1}{r_{3}}+\frac{1}{r_{4}})&\frac{j}{r_{4}}\\ 0&0&\frac{j}{r_{4}}&j(\frac{1}{r_{5}}-\frac{1}{r_{4}})\\ \end{pmatrix},\Omega^{-1}_{P}=\begin{pmatrix}0&0&-r_{3}&-r_{3}\\ 0&0&0&-r_{4}\\ r_{3}&0&0&0\\ r_{3}&r_{4}&0&0\\ \end{pmatrix}$.
Consider
$A_{1,1}:=\Omega_{P}^{-1}\big(\text{Hess}(J)_{P}+\text{Hess}(H_{t})_{P}\big)$
$=\begin{pmatrix}0&0&r_{5}-r_{4}-(2r_{3}+r_{4})t&\frac{r_{3}}{r_{5}}\big((r_{3}+2r_{4}-2r_{5})+(2r_{5}-r_{4})t\big)\\ 0&0&r_{3}+2r_{4}-2r_{5}+(r_{3}+r_{5})t&\frac{2(r_{4}-r_{5})^{2}+r_{3}r_{4}-r_{3}r_{5}-((r_{4}-r_{5})^{2}+r_{3}r_{5})t}{r_{5}}\\ r_{4}-r_{5}+r_{4}t&-r_{3}-r_{3}t&0&0\\ -r_{5}&-2r_{5}+r_{5}t&0&0\end{pmatrix}$
The quadratic polynomial corresponding to its characteristic polynomial is given by $\chi_{1,1}(X)=X^{2}+A(t)\,X+B(t)$, where
$A(t)$ $=[(r_{3}+r_{5})^{2}+2r_{4}j]t^{2}+2\!\left[r_{3}j+(r_{5}-r_{4})(r_{4}-2r_{5})\right]t+2r_{3}\,(r_{3}+3r_{4}-3r_{5})+5(r_{4}-r_{5})^{2}$
$B(t)$ $=j^{2}\,\bigl(-r_{4}t^{2}+r_{3}t+r_{4}t+r_{5}t+r_{3}+2r_{4}-2r_{5}\bigr)^{2}.$
By Definition 3.1, $P$ is of elliptic-elliptic type if $\chi_{1,1}(X)$ has two distinct negative roots, i.e. $A(t)>0,B(t)>0,\;A(t)^{2}>4B(t)$; and $P$ is of focus-focus type if $\chi_{1,1}(X)$ has no purely real roots, i.e. $A(t)^{2}<4B(t)$.
We first compare $A(t)^{2}$ with $4B(t)$.
$A(t)^{2}-4B(t)=\bigl((r_{3}+r_{5})t+2r_{3}+3r_{4}-3r_{5}\bigr)^{2}\,\bigl(at^{2}+bt+c\bigr),$
where $a=(r_{3}+r_{5})^{2}+4r_{4}j,b=-2\bigl(r_{3}r_{4}+r_{3}r_{5}+2r_{4}^{2}-3r_{4}r_{5}+r_{5}^{2}\bigr),c=(r_{4}-r_{5})^{2}$. Since $r_{3}>r_{5}$ and $r_{4}\geq\frac{1}{2}r_{5}$, we have $(r_{3}+r_{5})t+2r_{3}+3r_{4}-3r_{5}=(r_{3}+r_{5})t+2(r_{3}-r_{5})+3r_{4}-r_{5}>0$. Hence, $A(t)^{2}-4B(t)$ has the same sign as $f(t):=at^{2}+bt+c$. Since the discriminant $\Delta$ of $f(t)$ is $\Delta=b^{2}-4ac=16\,r_{3}r_{4}r_{5}j>0$ and $a>0$, $f(t)$ has two distinct roots; moreover, by the root formula, the two roots are $t^{\pm}$. We conclude that
  • when $t\in(-\infty,t^{-})\cup(t^{+},\infty)$, we have $A(t)^{2}>4B(t)$, so $\chi_{1,1}(X)$ has two distinct real roots.
  • when $t\in(t^{-},t^{+})$, we have $A(t)^{2}<4B(t)$, so $\chi_{1,1}(X)$ has no real roots, and it follows that $P$ is of focus-focus type.
Moreover, since $r_{3}>r_{5}>r_{4}$, we have $b=-2\big(r_{4}(r_{3}-r_{5})+r_{5}(r_{3}-r_{4})+r_{5}(r_{5}-r_{4})+2r_{4}^{2}\big)<0$ and $a+b=r_{3}^{2}-r_{5}^{2}+2r_{3}r_{4}+2r_{4}r_{5}>0$. Therefore, we know that $f(0)=c>0,f^{\prime}(0)=b<0,f(1)=a+b+c>0,f^{\prime}(1)=2a+b>0$, and hence the two roots $t^{\pm}$ of $f(t)$ must both be in $(0,1)$.
We now show that when $t\in[0,t^{-})\cup(t^{+},1]$, both roots of $\chi_{1,1}(X)$ are negative by showing that $A(t)>0,B(t)>0$. To check that $B(t)>0$, it suffices to show that $g(t):=-r_{4}t^{2}+(r_{3}+r_{4}+r_{5})t+(r_{3}+2r_{4}-2r_{5})$ has no roots in $[0,1]$. Notice that $g(0)=r_{3}-r_{5}+2r_{4}-r_{5}>0$, $g(1)=2r_{3}+2r_{4}-r_{5}>0$ and $g$ describes a parabola opening downwards, so $g(t)$ has no roots in $[0,1]$, i.e. $B(t)=j^{2}g(t)^{2}>0$ for all $t\in[0,1]$. As for $A(t)$, we notice that the coefficient of $t^{2}$ is positive and the discriminant of $A(t)$ is
 $4j\Bigl((r_{3}-r_{5})((r_{3}+r_{5})^{2}+7r_{3}r_{4}+5r_{4}^{2})+(2r_{4}-r_{5})(5r_{3}r_{4}+3r_{4}^{2})+7r_{4}r_{5}(r_{5}-r_{4})+3r_{4}^{3}\Bigr)<0$
Hence, $A(t)$ is always positive. Consequently, for $t\in[0,t^{-})\cup(t^{+},1]$, the singular point $P$ is of elliptic-elliptic type, and for $t\in(t^{-},t^{+})$, $P$ is of focus-focus type. ∎
6. Proof of Theorem 1.1
We are now ready to prove our main theorem.
Proof of Theorem 1.1.
We write $F_{t}=(J,H_{t})$ and $M=\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. By assumption, the moment images of $(J,\sqrt{H_{0}})$ and $(J,\sqrt{H_{1}})$ are both Delzant polygons. Hence, for $t=0,1$, the system is of toric type. Moreover, from the Delzant polygons (see Figure 1), $P$ is a maximum of $(H_{1})|_{J^{-1}(J(P))}$ and a minimum of $(H_{0})|_{J^{-1}(J(P))}$. Hence, it follows from Proposition 5.1 that $P$ is a transition point with transition time $t^{\pm}$.
We now show that for all $t\in[0,1]$, all other singular points of $F_{t}$ are non-degenerate of either elliptic-elliptic or elliptic-regular types. Fix a rank one point $m\in M$ that is not on the fixed surfaces $J^{-1}(|r_{1}-r_{2}|),J^{-1}(r_{1}+r_{2})$. By Lemma 2.5, the $S^{1}$-action generated by $J$ acts freely on $m$. By Proposition 4.2, if $c=J(m)\neq r_{3}+r_{4}-r_{5}$, then $[m]$ is an elliptic (in the Morse sense) singular point of $H_{t}^{\textrm{red},c}$ on $M^{\textrm{red},c}:=J^{-1}(c)/S^{1}$. Hence, by Lemma 3.7, $m$ is an elliptic-regular singularity of $F_{t}$. If $c=r_{3}+r_{4}-r_{5}$, then $M^{\textrm{red},c}$ is not a manifold. However, since $S^{1}$ acts freely on $m$, there is an $S^{1}$-invariant neighborhood around the orbit of $m$ on which $S^{1}$ acts freely. Hence, one can do symplectic reduction locally, and verify that $[m]$ is an elliptic singularity of $H_{t}^{\textrm{red},c}$ by computing its Hessian in the same way as in the proof of Proposition 4.2. Therefore, any rank one points that are not on the fixed surfaces are of elliptic-regular type. Finally, Proposition 3.9 implies that any rank one point on the fixed surfaces is of elliptic-regular type, and Proposition 3.9 and Proposition 4.2 together imply that any rank zero point on the fixed surfaces is of elliptic-elliptic type. Therefore, we conclude that for $t\in[0,1]\setminus\{t^{-},t^{+}\}$, $(M,\omega,F_{t})$ is a semitoric system, and when $t=t^{\pm}$, there are no degenerate singular points in $M\setminus\{P\}$. ∎